OpenStudy (anonymous):

If a^2=a+3, find a^3 in terms of a.

OpenStudy (anonymous):

multiply both sides by a: a^3 = a(a+3)

OpenStudy (anonymous):

a^2=a+3 a^2-a-3=0 Quadratic Formula!!!